Suppose a ruler balances at 37.1 cm with a total mass of 170.7 g as the counterweight. If the center of gravity of the ruler is at 50.6 cm, what is the mass of the ruler?

Express your answer to three significant figures and in units of grams
Answer #1

The ruler balances at 37.1 cm mark on the ruler.

The hanging mass (counter weight ) is M = 170.7 gm is at   mark on the ruler.

( you didn't give me procedure so i take 20 cm mark on the rular )

Center of gravity of ruler is at 50.6 cm mark ,

Mass of ruler is .

considering the torques about the point where the rules is balanced,

Mg*(37.1 - 20) = mg*(50.6 - 37.1 )

170.7*(37.1 - 20) = m *(50.6 - 37.1)

170.7*17.1 = m*13.5

m = 216.22 gm

Mass of the ruler m = 216.22 gm